# Practical privacy metrics for synthetic data

### v4

@@ Abstract
This paper explains how synthpop package 111see www.synthpop.org.uk  [^2] The new version 1.9-3 is currently available on Github at https://github.com/gillian-raab/synthpop has been extended to include functions to calculate measures of identity and attribute disclosure risk for synthetic data that measure risks for the records used to create the synthetic data. This paper is also appears as a vignette as pat of the versions of the package from 9.3 onwards. The basic function disclosure calculates identity disclosure for a set of quasi-identifiers (keys) and attribute disclosure for one variable specified as a target from the same set of keys. The second function multi.disclosure is a wrapper for the first and presents summary results for a set of targets. This short paper explains the measures of disclosure risk and documents how they are calculated. We recommend two measures: $RepU$ (replicated uniques) for identity disclosure and $DiSCO$ (Disclosive in Synthetic Correct Original) for attribute disclosure. Both are expressed a % of the original records and each can be compared to similar measures calculated from the original data. Experience with using the functions on real data found that some apparent disclosures could be identified as coming from relationships in the data that would be expected to be known to anyone familiar with its features. We flag cases when this seems to have occurred and provide means of excluding them.
@@ 1 Introduction
We expect that disclosure risks from SD to be lower than those from the original data. But in some cases, e.g. a large data set with only a few categorical variables, the disclosure risk from the original may already be low. To evaluate the protection from disclosure afforded by a synthesis method, the risks for SD must be compared with equivalent risks for the GT. We consider two types of disclosure risk:
  • identity disclosure: This refers to the ability to identify individuals in the data from a set of known characteristics that we will refer to as keys. Identity disclosure may be less relevant for completely synthesized [^3] Complete or full synthesis is when all values of all variables are replaced by synthetic values. This is in contrast to incomplete or partial synthesis where only some variables are replaced. data because there is no one-to-one correspondence between records in the original and SD. But it may still be of interest since it is an important factor for attribute disclosure.
  • attribute disclosure:This refers to the ability to find out from the keys something, not previously known, for an attribute associated with a record in the original data.
@@
The disclosure risk posed by SD can be reduced by using techniques from statistical disclosure control (sdc), such as aggregation of categories, smoothing of numeric values or removal of replicated uniques. These methods can be used to reduce disclosure risk by modifying the original data before synthesis, or the SD before its is released. Some such methods are already available in the synthpop package. These include categorising, top/bottom coding and smoothing for continuous variables, and the merging of small categories for factors. The removal of replicated uniques is another option available in synthpop.
@@ 2 A simple example
Here we illustrate the basic use of multi.disclosure. If the parameter targets is not specified, all the variables in the SD that are not part of keys are used as targets. The identity disclosure measures are UiO for original and repU for synthetic, and for attribute disclosure Dorig for original and DiSCO for synthetic. These and other measures will be explained in Sections 3.2 and 3.3.
First, a subset of 9 variables are selected from the SD2011 data (a survey on quality of life in Poland) that is available as part of the synthpop package. A single synthetic data set is created by the default method in synthpop: cart for each conditional distribution. Note that the variable income has values -8 that indicate not applicable, and the synthesis allows for this. The synthetic data object s1555an object of class synds has a component syn that is a single synthetic data set. The disclosure functions can also be used with synthetic data created by other methods either as single synthetic data sets or lists of repeated syntheses from the same original. Here we select 4 keys that represent items that might be known for members of this sample, or of the Polish population in 2011. By default,the 5 other variables become the targets: depress income ls marital (marital status), workab (intention to work abroad). The second target, income, is grouped with a target of 20 categories [^6] The algorithm attempts to create this number of groups with approximately equal sizes. Skew data often result in fewer groups. In this case 18 groups are formed from the numeric values, plus two groups from the -8 code and the missing values, giving a total of 20 We first use multi.disclosure to create a multi.disclosure object for these keys, and we print out its identity component.
$\begin{Schunk}$
$\begin{Sinput}$
@@ 3.1 Setting the scene
These disclosure measures are intended to assess what a person who only has access to the synthetic data can infer about known individuals who are present in the original data. We will use the term ”intruder” for such a person, though no malicious intent is implied. The intruder is assumed to have information for one or more individuals about the value of certain key variables that are present in the same format in the original and the synthetic data. They attempt first to see if the individual is present, and then to determine the value of other items in the data file that we refer to as targets. We are assuming a worst-case scenario where the intruder believes they are querying the original data. [^8] This may not be too unrealistic if the data are made available inadvertently, or if the intruder thinks that efforts to label the synthetic data as e.g. ”Fake Data” are thought just to be a cover up. It may also be a reqasonable measure to use without this scenario, since it compares the disclosiveness of the SD to that of the GT. Disclosure measures from the synthetic data are each compared to similar measures for someone with access to the original data. Here we will introduce the measures by an example. Formal definitions with notation and formulae are in Appendix 1. The first step in evaluating disclosure risk, as described here, is to identify a set of keys that might be expected to be known to an intruder. These keys are then combined to form a quasi-identifier that we designate as $q$. For example, if we have hospital records we might define age,sex date and hospital as keys and this would give a $q$ with levels such as ”78 | M | 1/1/2024 | WG” for a 78 year old man admitted to hospital WG on 1/1/2024.
@@ 3.2 Identity disclosure measures
The concept of k-anonymity is central to identity disclosure for microdata. First proposed in 1998 [20] it is discussed fully in [5]. A table is k-anonymous if a set of keys identifies at most $k$ individuals. Thus 2-anonymous data will never identify just one individual. Based on this idea, the percentage of records for which the keys identify just one individual give identity disclosure measures. Tables of $q$ values are produced from the synthetic and the original data. $UiO$ and $UiS$ are the percentages of records with keys where the table count is 1. An intruder checking out a record for their known set of keys will look for it in the synthetic data. Some records will not be in the SD and $UiOiS$ (Unique in Original in Synthetic) gives the % that would be found. These records are then checked for uniqueness in the synthetic data giving, $repU$ is the percentage of unique original records that are also unique in the SD.
The percentage $repU$ has been used as a disclosure measure to evaluate SD by [9] and by [17]  [^9] Jackson et al. in [9] argue that the denominator for $repU$ should be $N_{s}$ rather than $N_{d}$. This is inappropriate because our scenario is to consider the risk to the original data.. Replicated uniques are used in synthpop as part of the statistical disclosure control function, sdc, that includes the option of reducing disclosure risk by removing them from the SD. Nowok et al. [14] have evaluated this and give an example where this process has very little effect on utility. The function replicated.uniques [^10] For example by replicated.uniques(s2, ods, keys = c("sex","region","age","placesize")) also calculates $repU$ using a different method from the one described here.
@@ 3.3 Attribute disclosure measures
The attribute disclosure measure for the SD that is displayed in Figure 1 is our preferred measure, $DiSCO$ [16]. In section 3.3.1 we explain it via the example above. Other related disclosure measures, $DCAP$ and $TCAP$, have been proposed [6, 23, 22, 11, 10]. The disclosure function also computes these measures. Section 3.3.2 describes them and discusses circumstances where they differ from $DiSCO$.
The reason for our preference for $DiSCO$ compared to other proposed methods is that it is a mesure computed for all the records in the original data. The other measures thus do not include aspects of disclosure, for example that some records in the original not appearing in the SD, that are valid contributions to the disclosure reduction effected by the synthesis. The measure we describe as $DCAP$ is readily adapted to apply to all original records by replacing $N_{s}$ in its denominator with $N_{d}$. When the synthesis creates SD of the same size as the original $DCAP$ and $Dorig$ are identical. But, as we show in Figure 2, $DCAPorig$ assesses disclosure appropriately when synthetic data with fewer records than the original are created. This measure is added to the component attribattrib of the disclosure object created by the disclosure function.
The two measures $DiSCO$ and $DCAPorig$ can be considered as the extremities of a range of a more general disclosure measure when a record is considered disclosive if its probability of correct attribution reaches a threshold tautau. DiSCODiSCO corresponds to a tautau of 1.0 while $DCAPorig$ corresponds to a tautau of zero. This more general approach may be included in future releases of synthpopsynthpop.
Appendix 1 gives formal definitions of all disclosure measures discussed here.
3.3.1 Our preferred measure: Disclosive in Synthetic Correct in Original ($DiSCO$)
@@
  • Look up $q$ for each original record in the SD. The proportion found becomes $iS$ (in Synthetic).
  • Check if all records with the same $q$ have the same level of the target $t$. The proportion passing this further test becomes $DiS$ (Disclosive in Synthetic).
@@ 3.3.2 Other attribute disclosure measures
Previous work on attribute disclosure [6, 23] has used the Correct Attribution Probability ($DCAP$) as a disclosure measure for synthetic data. This is calculated as the average percentage of records correctly predicted from the synthetic data by a set of keys. Two different versions are described to deal with the case when some key combinations in the SD do not occur in the original, so that the attribution probability is undefined. In one case, that we will designate as $DCAP$, the undefined attribution probabilities are set to zero and the denominator is the total number of synthetic records. The second case $DCAP_{alt}$ [^11] This approach is sometimes referred to as $TCAP_{undefined}$ [23]. uses the number of synthetic records present in the synthetic data as the denominator, so that $DCAP_{alt}\geq DCAP$.
@@
The related quantity $TCAP$ restricts the probabilities to those over a certain limit, and most commonly this limit is $1.0$ making the numerator of the $TCAP$ measures identical to that for $DiSCO$. One possibility for a denominator would be to use just the number of synthetic records with key combinations in the original, as for $DCAP_{alt}$ [^12] This measure was what was reported as $TCAP$ in Versions of $Synthpop$ before Version 1.9.3. But evaluations using $TCAP$ [11, 10] use an even smaller denominator; the number of synthetic records with keys present in the original and where the target probaility is 1.0. This can be described as the proportion correctly predicted for all apparently disclosive records; we refer to this measure as $TCAP$ and note that its smaller denominator implies that $TCAP\geq TCAP_{alt}\geq DiSCO$ when the synthetic data set(s) are the same size as the original.
@@
All of these measures are shown in Table 1 for 3 of the targets in the synthesis of the 5,000 records from the example above. We can see that the inequalities above all hold. For this case the ordering of the disclosive risk by target is consistent for all measures. For the three measures that share the same numerator ($DiSCO$, $TCAP_{alt}$ and $DCAP$) we see that the reduced denominators have a large effect in this example, so that $TCAP$ is around three times $DiSCO$.
DiSCO DCAP DCAPalt TCAPalt TCAP
ls 12.63 21.97 29.57 17.00 36.71
depress 9.21 16.23 21.85 12.40 28.57
income 6.16 11.34 15.27 8.29 19.96
@@
Table 1: Different disclosure measures (%) for 3 targets from the example above when number of synthetic records equals the number of original records (5000).
@@
Another way thatt he disclosure risk to the original records can be reduced is to release synthetic data sets that are smaller than the original. This is the approach that has been used when UK Census data has been released [2] or for low-fidelity Census data from Scotland [^13] https://www.researchdata.scot/news-and-insights/synthetic-census-data-now-available-for-research/ accessed 7/5/25. Figure 2 illustrates how the different disclosure measures assess the risks in this example for sizes of SD from 5000, as the original, to only 250 records. Of the measures in Table 1, only $DiSCO$ and $DCAPorig$ adjust the attribute disclosure risk for the reduced number of records in the SD. The $TCAP$ measure is almost independent of the size of the SD. The other measures all give misleadingly high disclosure risks for small samples.
Figure 2: Plot of disclosure measures for synthetic data of size ($k$) from 250 to 5000.
Figure 2 displays

### v5

@@ Abstract
This paper explains how synthpop package 111see www.synthpop.org.uk has been extended to include functions to calculate measures of identity and attribute disclosure risk for synthetic data that measure risks for the records used to create the synthetic data. The basic function disclosure calculates identity disclosure for a set of quasi-identifiers (keys) and attribute disclosure for one variable specified as a target from the same set of keys. The second function multi.disclosure is a wrapper for the first and presents summary results for a set of targets. This short paper explains the measures of disclosure risk and documents how they are calculated. We recommend two measures: $RepU$ (replicated uniques) for identity disclosure and $DiSCO$ (Disclosive in Synthetic Correct Original) for attribute disclosure. Both are expressed a % of the number of original records and each can be compared to similar measures calculated from the original data on its own. Experience with using the functions on real data found that some apparent disclosures could be identified as coming from relationships in the data that would be expected to be known to anyone familiar with its features. We flag cases when this seems to have occurred and provide means of excluding them. This paper was written as a vignette for the R package synthpop version 1.9-3.
@@ 1 Introduction
We expect that disclosure risks from SD to be lower than those from the original data. But in some cases, e.g. a large data set with only a few categorical variables, the disclosure risk from the original may already be low. To evaluate the protection from disclosure afforded by a synthesis method, the risks for SD must be compared with equivalent risks for the original ground truth data (GT). We consider two types of disclosure risk:
  • identity disclosure: This refers to the ability to identify individuals in the data from a set of known characteristics that we will refer to as keys. Identity disclosure may be less relevant for completely synthesized [^2] Complete or full synthesis is when all values of all variables are replaced by synthetic values. This is in contrast to incomplete or partial synthesis where only some variables are replaced. data because there is no one-to-one correspondence between records in the GT and the SD. But it may still be of interest since it is an important factor for attribute disclosure.
  • attribute disclosure:This refers to the ability to find out from the keys something, not previously known, for an attribute associated with a record in the GT.
@@
The disclosure risk posed by SD can be reduced by using techniques from statistical disclosure control (sdc), such as aggregation of categories, smoothing of numeric values or removal of replicated uniques. These methods can be used to reduce disclosure risk by modifying the GT before synthesis, or the SD before its is released. Some such methods are already available in the synthpop package. These include categorising, top/bottom coding and smoothing for continuous variables, and the merging of small categories for factors. The removal of replicated uniques is another option available in synthpop.
@@
The Appendix introduces the notation we will use to define the disclosure measures that can be computed by the disclosure functions in the synthpopsynthpop package. Our approach involves creating a composite variable, $q$, from the quasi-identifiers (keys) for the original and synthetic data. Identity disclosure measures are calculated from univariate tables of $q$ for the GT and the SD. Attribute disclosure measures for a target variable $t$ are obtained by tabulating $t$ against $q$ for the GT and the SD. For our initial example in Section 2 and most of the later development we assume thet the number of records created in the SD ($N_{s}$) is the same as the number in the GT ($N_{d}$). The exception is Section 3.3.2 where we show how different proposed measures for attribute disclosure are affected by reducing the number of synthetic records ($N_{s}$) that are released.
@@ 2 A simple example
Here we illustrate the basic use of multi.disclosure. If the parameter targets is not specified, all the variables in the SD that are not part of keys are used as targets. The identity disclosure measures are UiO for the original and repU for synthetic, and for attribute disclosure Dorig for original and DiSCO for synthetic. These and other measures will be explained in Sections 3.2 and 3.3.
First, a subset of 9 variables are selected from the SD2011 data (a survey on quality of life in Poland) that is available as part of the synthpop package. A single synthetic data set is created by the default method in synthpop: cart for each conditional distribution. Note that the variable income has values -8 that indicate not applicable, and the synthesis allows for this. The synthetic data object s1444an object of class synds has a component syn that is a single synthetic data set. The disclosure functions can also be used with synthetic data created by other methods either as single synthetic data sets or lists of repeated syntheses from the same original. Here we select 4 keys that represent items that might be known for members of this sample, or of the Polish population in 2011. By default,the 5 other variables become the targets: depress income ls marital (marital status), workab (intention to work abroad). The second target, income, is grouped with a target of 20 categories [^5] The algorithm attempts to create this number of groups with approximately equal sizes. Skew data often result in fewer groups. In this case 18 groups are formed from the numeric values, plus two groups from the -8 code and the missing values, giving a total of 20 We first use multi.disclosure to create a multi.disclosure object for these keys, and we print out its identity component.
@@ 3.1 Setting the scene
These disclosure measures are intended to assess what a person who only has access to the synthetic data can infer about known individuals who are present in the original data. We will use the term ”intruder” for such a person, though no malicious intent is implied. The intruder is assumed to have information for one or more individuals about the value of certain key variables that are present in the same format in the original and the synthetic data. They attempt first to see if the individual is present, and then to determine the value of other items in the data file that we refer to as targets. We are assuming a worst-case scenario where the intruder believes they are querying the original data. [^7] This may not be too unrealistic if the data are made available inadvertently, or if the intruder thinks that efforts to label the synthetic data as e.g. ”Fake Data” are thought just to be a cover up. It may also be a reqasonable measure to use without this scenario, since it compares the disclosiveness of the SD to that of the GT. Disclosure measures from the synthetic data are each compared to similar measures for someone with access to the original data. Here we will introduce the measures by an example. Formal definitions with notation and formulae are in the Appendix. The first step in evaluating disclosure risk, as described here, is to identify a set of keys that might be expected to be known to an intruder. These keys are then combined to form a quasi-identifier that we designate as $q$. For example, if we have hospital records we might define age,sex date and hospital as keys and this would give a $q$ with levels such as ”78 | M | 1/1/2024 | WG” for a 78 year old man admitted to hospital WG on 1/1/2024.
@@ 3.2 Identity disclosure measures
The concept of k-anonymity is central to identity disclosure for microdata. First proposed in 1998 [20] it is discussed fully in [5]. A table is k-anonymous if a set of keys identifies at most $k$ individuals. Thus 2-anonymous data will never identify just one individual. Based on this idea, the percentage of records for which the keys identify just one individual give identity disclosure measures. Tables of $q$ values are produced from the SD and the GT. $UiO$ and $UiS$ are the percentages of records with keys where the table count is 1. An intruder checking out a record for their known set of keys will look for it in the synthetic data. Some records will not be in the SD and $UiOiS$ (Unique in Original in Synthetic) gives the % that would be found. These records are then checked for uniqueness in the synthetic data giving, $repU$ is the percentage of unique records in GT that are also unique in the SD.
The percentage $repU$ has been used as a disclosure measure to evaluate SD by [9] and by [17]  [^8] Jackson et al. in [9] argue that the denominator for $repU$ should be $N_{s}$ rather than $N_{d}$. This is inappropriate because our scenario is to consider the risk to the original data, as we discuss for attribute disclosure in Section 3.2.2.. Replicated uniques are used in synthpop as part of the statistical disclosure control function, sdc, that includes the option of reducing disclosure risk by removing them from the SD. Nowok et al. [14] have evaluated this and give an example where this process has very little effect on utility. The function replicated.uniques [^9] For example by replicated.uniques(s2, ods, keys = c("sex","region","age","placesize")) also calculates $repU$ using a different method from the one described here.
@@ 3.3 Attribute disclosure measures
The attribute disclosure measures displayed in Figure 1 are our preferred measures, $DiSCO$, for the SD and $Dorig$ for the GT [16]. In section 3.3.1 we explain them via the example above. Other related disclosure measures, $DCAP$ and $TCAP$ and their variants, have been proposed . We prefer $DiSCO$ over $TCAP$ and its variants [11, 10] because it measures all the ways synthesis is effective in reducing attribute disclosure. A further measure $DCAP_{d}$, a modification of [6, 22], also measures all these aspects. The measures $DiSCO$ and $DCAP_{d}$ are included in the attribattrib component of the output from the disclosure function. For completeness, all the other measures are computed by disclosuredisclosure and returned in the $allCAPs$ component of the disclosure object.
The two measures $DiSCO$ and $DCAP_{d}$ are the only two measures that show how the disclosure risk is reduced by releasing SD with fer records than GT, see Figure2. They can be considered as the extremities of a range of a more general disclosure measure when a record is considered disclosive if its probability of correct attribution reaches a threshold tautau. DiSCODiSCO corresponds to a tautau of 1.0 while $DCAP_{d}$ corresponds to a tautau of zero. This more general approach may be included in future releases of synthpopsynthpop. Section 3.3.2 discusses all these other measures and formal definitions of all measures are in the Appendix.
3.3.1 Disclosive in Synthetic Correct in Original: DiSCODiSCO
@@
  • Look up $q$ in the SD for each original record in the GT. The proportion found becomes $iS$ (in Synthetic).
  • Check if all records with the same $q$ have the same level of the target $t$. The proportion of all GT records passing this further test becomes $DiS$ (Disclosive in Synthetic).
@@ 3.3.2 Other attribute disclosure measures
Previous work on attribute disclosure [6, 22] has used the Correct Attribution Probability ($CAP$) as a disclosure measure for synthetic data. This is the proportion for a record in the SD where its $q$ correctly predicts the attribute level of the target, $t$, in the original. Two versions are described, each with the total $CAP$ in the numerator, but different denominators. Some combinations of the keys, $q$, found in the SD may not exist in the original, so that their $CAP$ is undefined. One version, that we will designate as $DCAP_{b}$, ignores the records with undefined probabilities and uses as a denominator $N_{b}$, the number of synthetic records where the value of $q$ occurs in the original. [^10] This approach is sometimes referred to as $DCAP_{undefined}$ [22]. A second version ($DCAP_{s}$) sets the undefined attribution probabilities to zero for records where $CAP$ is undefined and uses as denominator ($N_{s}$) the total number of synthetic records. Note that $DCAP_{b}\geq DCAP_{s}$ because $N_{b}\leq N_{s}$. If the total mumber of original records is used as the dsenominator we get the measure $DCAP_{d}$. This is the same as $DCAP_{s}$ when the SD has the same number of records as the GT, but will be lower when smaller SD samples are generated.
@@
The related quantity $TCAP$, and its variants. restricts the probabilities to those over a certain limit, and most commonly this limit is $1.0$. WIth this choice the numerator of the $TCAP$ measures identical to that for $DiSCO$, the proportion of synthetic records correctly predicted with certainty from the SD. The same two choices of denominator as were used for $DCAP$ could be used. One choice $TCAP_{s}$ would be to use the denominator, $N_{s}$ and the other $TCAP_{b}$ would be to use the denominator, $N_{b}$. Note that when $N_{s}=N_{d}$, the measure $TCAP_{s}$ will be identical to $DiSCO$. However the $TCAP$ that is defined in [11, 10] does not use either of these denominators, but uses as a denominator $N_{bp}$ the number of synthetic records with $q$ present in the original and where any level of the target $t$ is predicted with certainty, but not necessarily correct. Another way of describing this $TCAP$ is as the proportion of apparently disclosive recordsin the SD that are correct in the original. The numerators and denominators used for each of these measures are summarised in Table 1.
Numerator
Sum of proportions correct Total number predicted with certainty
Denominator CAP CAP = 1
$N_{d}$ $DCAP_{d}$ DiSCODiSCO
$N_{s}$ $DCAP_{s}$ $TCAP_{s}$
$N_{b}$ $DCAP_{b}$ $TCAP_{b}$
$N_{bp}$ TCAPTCAP
Table 1: Summary of numerators and denominators for $DCAP$ and TCAPTCAP measures.
@@
All of these measures are shown in Table 2 for 3 of the targets in the synthesis of the 5,000 records from the example above. We can see that the inequalities above all hold. For this case the ordering of the disclosive risk by target is consistent for all measures. Because the number of records in the GT and the SD are the same, $DCAP_{s}$ is identical to $DCAP_{d}$ and $TCAP_{s}$ is identical to $DiSCO$. For measures that share the same numerator, Table 1, we see that the reduced denominators have a large effect in this example, so that $TCAP$ is around three times $DiSCO$.
$DCAP_{s}$ and $DCAP_{d}$ $DCAP_{b}$ $TCAP_{d}$ and DiSCODiSCO $_{b} TCAPTCAP$
ls 21.97 29.57 12.63 17.00 36.71
depress 16.23 21.85 9.21 12.40 28.57
income 11.34 15.27 6.16 8.29 19.96
@@
Table 2: DCAP and TCAP variants (%) for 3 targets from the example above when number of synthetic records equals the number of original records (5000). Average results for 5 SDs.
@@
For all the examples we have examined so far SD has been generated with sample size $N_{s}$ the same as the number of original records $N_{d}$; $N_{s}=N_{d}$. Another way that the disclosure risk to the original records can be reduced is to release synthetic data sets that are smaller than the original. This is the approach that has been used when UK Census data has been released [2] or for low-fidelity Census data from Scotland [^11] https://www.researchdata.scot/news-and-insights/synthetic-census-data-now-available-for-research/ accessed 7/5/25.
Figure 2: Plot of disclosure measures for 3 targets for synthetic data of size ($k$) from 250 to 5000. Note that the legends apply to all figures, with type of line indicating the numerators and colours the denominators of each measure.
Figure 2 illustrates how the different disclosure measures assess the risks in this example for sizes of SD from 5000, as the original, to only 250 records. Of the measures in Table 1, only $DiSCO$ and $DCAP_{d}$ adjust the attribute disclosure risk correctly for the reduced number of records in the SD. The $TCAP$ measure is almost independent of the size of the SD. The other measures all give misleadingly high disclosure risks for small samples.We can see how the pairs of measures ($DiSCO$, $TCAP_{s}$) and ($DCAP_{d}$$DCAP_{s}$) diverge for $N_{s}$ values lower than $N_{d}$.
Referring back to section 3.3.1 we can see how these measures relate to the three bullet points giving the stages by which an intruder takes to identify an attribute disclosure. The measure $TCAP$ assesses the intruder’s success rate only at the final bullet point, starting with records in $DiS$. $TCAP_{s}$ measures the success of the last two steps, from $iS$, but only $DiSCO$ includes the disclosure risk from all 3 steps.